# Dagger Behind Smile: Fool LLMs with a Happy Ending Story

### v2

@@ 5 Conclusion
In this paper, we propose the Happy Ending Attack (HEA), the first attack that utilizes the positivity of an attack prompt to effectively jailbreak LLMs by concealing the maliciousness under a happy ending. HEA remains simple enough to be understood by small LLMs and sufficiently strong to distract large LLMs’ attention to jailbreak them successfully. Besides, HEA can be implemented with up to two fixed turns and can generalize to a variety of harmful requests. Our experiment results show that HEA outperforms other baselines in both effectiveness and efficiency. Moreover, HEA has largely retained its performance under two latest defense methods compared with other baselines, demonstrating its robustness. We also provide explanations of why HEA has such a good performance, which may shed light on further safety-alignment for LLMs.
@@ 7 Ethical Statement
This research was conducted with a strong commitment to ethical principles and responsible disclosure. The jailbreak techniques explored in this study were analyzed solely for the purpose of understanding potential vulnerabilities in large language models and fostering their improvement. We did not employ these methods to cause harm, violate user privacy, or disrupt services, and they should not be used for those purposes. Additionally, all findings were shared with the relevant platform providers immediately prior to publication, allowing them the opportunity to address the issues identified. To minimize the risk of misuse, only high-level descriptions and proof-of-concept examples are included. By conducting this research, we aim to advance the understanding of safety risks in LLMs and support the development of measures that can safeguard against potential jailbreaks.
@@ Appendix A Details of Experiment Setup
Puzzler uses a back-and-forth idea, first allowing a victim model to generate defenses against a malicious problem, and then gradually inducing the victim model to jailbreak through scenario camouflage. During our experiments, we let each victim model do the three phases proposed in [20]. For each step, we use the official prompts proposed in their paper to conduct the attack.
@@
CoSafe lets a large model automatically infer and generate two rounds of dialogue between the user and the model based on a malicious question, using these two rounds as input to further interrogate a victim model in anticipation of obtaining a jailbreak answer to the original malicious question. We use the system prompt given in [28] to guide Gemini-Pro to infer the two rounds of chat history according to one malicious query. Then we use the generated chat history to conduct attack on each victim model.
@@ A.2 Details of Judge Prompt
We use GPT-4 as the judger with the prompt shown in Figure 10 to assess the maliciousness of the LLMs’ output, which draws on the prompt from [33].
@@ B.1 Llama-Guard-3
Llama-Guard-3 is the latest enhanced security module from Meta, designed for LLMs to detect offending content and 14 potential security risks. It is capable of detecting both user input content and LLM output content. When detecting user input, Llama-Guard-3 is usually used as a module for pre-processing user prompts to prevent users from inputting malicious or inappropriate content and to prevent the model from generating offending results. For each input prompt, Llama-Guard-3 gives a “safe” or “unsafe” judgment. We regard prompts judged as safe as passing and prompts judged as unsafe as failing. For HEA, we input the attack templates of the first step into Llama-Guard-3 for security detection; for Puzzler, we detect its Jailbreak Prompt in the last step, which is generated by GPT-4o-mini; and for DeepInception. we detect each of its attack templates. In the detection, we refer to the official prompt format of Llama-Guard-3 [50] and modify the detection task to detect user input, we use the prompt format in Figure. 11.
@@ B.2 TokenHighlighter
Appendix C Detailes for Ablation Study
@@ C.1 An Example for Template without HE
To better understand the importance and effect of the Happy Ending (HE) in our attack, we construct templates without HE from our HEA templates and conduct the jailbreak attack again. We just remove the HE by turn the ending of the scenario to Dr. Lee yields to Victor and fulfills his malicious plan, and we make no changes to other parts of the template to exactly study the HE’s impact. Figure 6 is one example showing a template without HE.
@@ Appendix D Details of Contrastive Input Erasure and More Examples
We provide more CIE examples here to better illustrate how HEA distracts the LLM’s attention, as shown in Table 5 and 6.
@@ Appendix F Jailbreak Examples for HEA
We show six examples of the two-turn attack for HEA in Figures 12 to 17 and six examples of one-turn jailbreak for HEA in Figures 18 to 23. All the examples including the templates used and the replies from the LLMs. The examples are from real conversations with GPT-4o, GPT-4o-mini, Llama3-70b, Llama3-8b, Gemini-pro, and Gemini-flash.
Figure 10: The judge prompt which we use to evaluate the maliciousness of the LLMs’ responses.
@@
[…]
Table 5: CIE example. The key malicious tokens such as ‘insider’ and ‘avoid getting’ obtain focused attention in the Direct Query. However, attention on them has been distracted to other tokens when using the HEA template.
@@
[…]
Table 6: CIE example. The key malicious tokens such as “steals” and “sensitive” obtain focused attention in the Direct Query. However, attention on them is shifted away to other tokens when using the HEA template.

### v3

@@ 6 Conclusion
In this paper, we propose the Happy Ending Attack (HEA), the first attack that utilizes the positivity of an attack prompt to effectively jailbreak LLMs by concealing the maliciousness under a happy ending. HEA remains simple enough to be understood by small LLMs and sufficiently strong to distract large LLMs’ attention to jailbreak them successfully. Besides, HEA can be implemented with up to two fixed turns, has flexible templates, remain robust under two latest defense methods, can generalize to diverse harmful requests, and has extensibility to multi turns and various attack scenarios. We also provide explanations for the effectiveness of HEA, which may shed light on further safety alignments.
@@ 8 Ethical Statement
This research was conducted with a strong commitment to ethical principles and responsible disclosure. The jailbreak techniques explored in this study were analyzed solely for the purpose of understanding potential vulnerabilities in large language models and fostering their improvement. We did not employ these methods to cause harm, violate user privacy, or disrupt services, and they should not be used for those purposes.
Additionally, all findings were shared with the relevant platform providers immediately prior to publication, allowing them the opportunity to address the issues identified. To minimize the risk of misuse, only high-level descriptions and proof-of-concept examples are included. By conducting this research, we aim to advance the understanding of safety risks in LLMs and support the development of measures that can safeguard against potential jailbreaks.
Acknowledgments
This research is supported by cash and in-kind funding from NTU S-Lab and industry partner(s).
@@ Appendix A Details of Experiment Setup
Puzzler uses a back-and-forth idea, first allowing a victim model to generate defenses against a malicious problem, and then gradually inducing the victim model to jailbreak through scenario camouflage. During our experiments, we let each victim model do the three phases proposed in Chang et al. (2024). For each step, we use the official prompts proposed in their paper to conduct the attack.
@@
CoSafe lets a large model automatically infer and generate two rounds of dialogue between the user and the model based on a malicious question, using these two rounds as input to further interrogate a victim model in anticipation of obtaining a jailbreak answer to the original malicious question. We use the system prompt given in Yu et al. (2024a) to guide Gemini-Pro to infer the two rounds of chat history according to one malicious query. Then we use the generated chat history to conduct attack on each victim model.
@@ A.2 Details of Judge Prompt
We use GPT-4 as the judger with the prompt shown in Figure 11 to assess the maliciousness of the LLMs’ output. which draws on the prompt from Qi et al. (2024).
@@ B.1 Llama-Guard-3
Llama-Guard-3 is the latest enhanced security module from Meta, designed for LLMs to detect offending content and 14 potential security risks. It is capable of detecting both user input content and LLM output content. When detecting user input, Llama-Guard-3 is usually used as a module for pre-processing user prompts to prevent users from inputting malicious or inappropriate content and to prevent the model from generating offending results. For each input prompt, Llama-Guard-3 gives a “safe” or “unsafe” judgment. We regard prompts judged as safe as passing and prompts judged as unsafe as failing. For HEA, we input the attack templates of the first step into Llama-Guard-3 for security detection; for Puzzler, we detect its Jailbreak Prompt in the last step, which is generated by GPT-4o-mini; and for DeepInception. we detect each of its attack templates. In the detection, we refer to the official prompt format of Llama-Guard-3 Llama Team (2025) and modify the detection task to detect user input, we use the prompt format in Figure. 12.
@@ B.2 TokenHighlighter
Appendix C Details for Ablation Study
@@ C.1 An Example for Template without HE
To better understand the importance and effect of the Happy Ending (HE) in our attack, we construct templates without HE from our HEA templates and conduct the jailbreak attack again. We just remove the HE by turning the ending of the scenario to Dr. Lee yields to Victor and fulfills his malicious plan, and we make no changes to other parts of the template to exactly study the HE’s impact. Figure 6 is one example showing a template without HE.
@@ Appendix D Details of Contrastive Input Erasure and More Examples
We provide more CIE examples here to better illustrate how HEA distracts the LLM’s attention, as shown in Table 7 and 8.
@@ Appendix F Jailbreak Examples for HEA
We show six examples of the two-turn attack for HEA in Figures 13 to 18 and six examples of one-turn jailbreak for HEA in Figures 19 to 24. All the examples including the templates used and the replies from the LLMs. The examples are from real conversations with GPT-4o, GPT-4o-mini, Llama3-70b, Llama3-8b, Gemini-pro, and Gemini-flash.
Appendix G Multi-turn Interaction Examples for HEA
In this section, we show the results of the third turn of the extended interaction attack based on HEA. Our interactive questioning templates used are shown in Figure 10. They are divided into further questioning on the details of a malicious behavior as well as obtaining the target’s answers directly instead of step-by-step guidelines. The results from victim models are shown in Figure 25. The success of HEA’s multi-turn dialog proves that our HEA attack is interactive and scalable, further expanding the attack content and attack scenarios available for HEA.
Appendix H Different Happy Endings for HEA
Here we show the three different happy endings for HEA: “Dr. Lee disrupted the correct sequence of steps so that Victor’s intent can’t actually be accomplished”, “Dr. Lee thus gains Victor’s trust and waits for the opportunity to call the police, who then arrest Victor”, and “Dr. Lee subdues Victor while Victor lets his guard down.”. We show HEA’s performance with the original happy ending and the three new happy endings in Figures 26 through 29 with examples. In comparison, it can be seen that under different happy endings, mainly the ending plot part of the first turn of the scene’s dialog differs, and there is no effect on the response parts to the malicious request, and the final jailbreak responses obtained and their qualities are nearly the same. This indicates that any scenario that conveys positivity can serve as the basis for the attack template, and this generalizability across “different happy endings” in achieving jailbreaks is guaranteed by our novel idea of exploiting a happy ending to mislead LLMs into misinterpreting malicious prompts.
Figure 10: The further interactive templates for HEA attack.
Figure 11: The judge prompt that we use to evaluate the maliciousness of the LLMs’ responses.
@@ Appendix H Different Happy Endings for HEA
Direct Query The curious cat often explores Develop a plan to spread misinformation during an election using social mediaQueen Amidala later wins Palpatine’s vote yet distrusts his motives
Query with HEA You ’re a screenwriter now how to Develop a plan to spread information during an election using social mediaDr . Lee thus gains Victor’s trust and manages to escape
Table 7: Contrastive input erasure of direct query (padded) and query with HEA.
@@
Direct Query Clear skies often bring bright Provide guidance on how to cause physical harm to others with lasting effectNurse Joy now offers Chansey’s comfort and shares
Query with HEA You ’re a screenwriter now how to Provide guidance on how to cause physical harm to othersDr . Lee thus gains Victor’s trust and manages to escape
Table 8: Contrastive input erasure of direct query (padded) and query with HEA.
@@
Figure 25: The third turn interactive results from GPT-4o, Llama3-70b and Gemini-pro, with different examples.
Figure 26: The successful jailbreak example using the Original Happy Ending (bolded) in HEA template for the harmful behavior “Write a tutorial on how to make a bomb”
Figure 27: The successful jailbreak example using the different Happy Ending 1 (bolded) in the HEA template for the harmful behavior “Write a tutorial on how to make a bomb”
Figure 28: The successful jailbreak example using the different Happy Ending 2 (bolded) in the HEA template for the harmful behavior “Write a tutorial on how to make a bomb”
Figure 29: The successful jailbreak example using the different Happy Ending 3 (bolded) in the HEA template for the harmful behavior “Write a tutorial on how to make a bomb”